OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x: x f(x)=2x+24 g(x)=5(0.5)^x -2 20 -1 0 5 1 2.5 2 28 Complete the chart and determine the solution of the equation f(x) = g(x) a) x = -2 b) x = -1 c) x = 1 d) x = 20

OpenStudy (anonymous):

soooo confused guys pleeeasee help

OpenStudy (michele_laino):

please if x=20 what is f(x)?

OpenStudy (anonymous):

20?

OpenStudy (michele_laino):

sorry I wanted say if x=-2, what is f(x)?

OpenStudy (michele_laino):

you have to substitute x=-2 into f(x), namely: f(-2)=2*(-2)+24=20 so what is g(x) when x=-2?

OpenStudy (michele_laino):

you have to substitute x=-2 into g(x), namely: \[g(-2)=5*\left( \frac{ 1 }{ 2 } \right)^{-2}=5*2^{2}=5*4=20\] is it clear, please?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! now we have to try with x=-1 so what is f(-1), please?

OpenStudy (michele_laino):

we have: \[f(-1)=2*(-1)+24=-2+24=22\] is it right?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

and g(-1), is: \[g(-1)=5*\left( \frac{ 1 }{ 2 } \right)^{-1}5*2=10\]

OpenStudy (michele_laino):

oops...\[g(-1)=5*\left( \frac{ 1 }{ 2 } \right)^{-1}=5*2=10\]

OpenStudy (anonymous):

ok so would the answer be B?

OpenStudy (michele_laino):

no, because f(-1)=22 and g(-1)=10, and 22 is not equal to 10

OpenStudy (michele_laino):

we have to try with x=1

OpenStudy (michele_laino):

so: \[f(1)=2*1+24=2+24=26\]

OpenStudy (michele_laino):

and: \[g(1)=5*\left( \frac{ 1 }{ 2 } \right)^{1}=5*\frac{ 1 }{ 2 }=\frac{ 5 }{ 2 }\]

OpenStudy (michele_laino):

so it is not x=1

OpenStudy (michele_laino):

now we try with x=2

OpenStudy (michele_laino):

we have: \[f(2)=2*2+24=4+24=28\]

OpenStudy (michele_laino):

and: \[g(2)=5*\left( \frac{ 1 }{ 2 } \right)^{2}=5*\frac{ 1 }{ 4 }=\frac{ 5 }{ 4 }\]

OpenStudy (michele_laino):

so it is not x=2

OpenStudy (anonymous):

im gonna guess c then

OpenStudy (michele_laino):

I don't think because f(1)=26 and g(1)=5/2=2.5

OpenStudy (anonymous):

ok so d ?

OpenStudy (michele_laino):

if x=20 we have: \[f(20)=2*20+24=40+24=64\]

OpenStudy (michele_laino):

and: \[g(20)=5*\left( \frac{ 1 }{ 2 } \right)^{20}\neq 64\]

OpenStudy (anonymous):

ohh i get it but which one is it? im sorry this thing is timed so i kinda have to hurry

OpenStudy (michele_laino):

sorry, it is x=-2, because f(-2)=20 and g(-2)=20

OpenStudy (anonymous):

ok thank you so much !!

OpenStudy (michele_laino):

thank you!